### v1

@@ Appendix B Additional data details - Tables & Figures
This appendix documents the sample composition and key distributional facts that underlie the main analyses. We report continuous/count variables in Table B.1, binary indicators in Table B.3, and multi–category variables in Table B.3. Figure B.1 summarizes how the mass of contributions moves across ranges over the ten rounds, and Table B.4 provides corresponding headcounts by round and type. All the percentages are balanced and indicative to the real percentages of the population in Honduran Villaes (see Oles2024Maternal; airoldi2024induction; papamichalis2025educational for comparison).
@@ Appendix E IV theory & diagnostics
Each round $t$, player $i$ in a group of size $N=5$ chooses $c_{it}\in[0,12]$. Instantaneous utility combines material payoffs, altruism, and norm aversion:
$\pi_{it}=\frac{b}{N}\!\left(c_{it}+(N-1)\bar{c}_{-i,t}\right)-\kappa c_{it}+d_{i}c_{it}^{\alpha}-h_{i}\exp\!\left\{k\big(c_{it}-\bar{c}_{-i,t-1}\big)^{2}\right\},$
with $\alpha\in(0,1)$, $d_{i}>0$, $h_{i}\in[0,1]$, $k>0$, and $\bar{c}_{-i,t}$ the leave–one–out peer mean. Under $b/N<\kappa<b$, best–reply dynamics admit two locally attracting fixed points, generating the two–basin structure we document.
@@
$Z^{SS}_{i,t-1}\;=\;\sum_{k}s_{g,-i}^{(k)}\,\mu_{t-1}^{(k)},\qquad\text{and its leave–one–village variant}\quad Z^{LOV}_{i,t-1}\;=\;\sum_{k}s_{g,-i}^{(k)}\,\mu_{t-1}^{(-v),(k)},$
where $s_{g,-i}^{(k)}$ is the LOO group share for trait $k$ and $\mu_{t-1}^{(-v),(k)}$ is the cross–village mean at t1t-1 excluding village $v$. Random assignment within villages and the predetermined nature of traits support relevance and exclusion. Because avatars are anonymous, composition can affect $c_{it}$ only via peers’ contributions.
@@ Appendix F Structural back‑out
We summarize behavior using two primitives. First, an altruism weight $d_{i}$ governing the intrinsic marginal benefit from giving (warm–glow). Second, an effective norm–pull $\phi_{i}$ that penalizes deviations from the group’s lagged mean. In the first–order condition, only the product of a curvature parameter and a conformity scale is identified, so we collect these in
$\phi_{i}\equiv 2k\,h_{i},$
and estimate $(d_{i},\phi_{i})$ by minimum distance on each player’s within–session choices using the leave–one–out lagged peer mean (units: Lempiras). [^12] All peer means exclude the focal player. Variables are in lempiras (0–12). Identification is partial: only the product $\phi_{i}=2kh_{i}$ is point–identified from the quadratic deviation cost; the mapping to $(k,h_{i})$ is conventional and used for scaling only. We fix $\alpha=\tfrac{1}{2}$ in the warm–glow term and show sensitivity to $\alpha\in\{0.3,0.7\}$ in SI. Data handling follows the main text: rounds 221010 for lags; a small ridge ($10^{-3}$) when $c_{it}=0$; and 99th–percentile winsorization for the back–out residual.
Three facts emerge. (a) The distribution of $d_{i}$ is broad (median $\approx 2.8$; IQR 1.91.93.43.4); plugging these into the individual best–reply delivers an implied high fixed point at the 12 L cap for essentially all players, i.e., altruism is in principle sufficient to support high contributions. (b) The inferred norm–pull $\phi_{i}$ is near zero for the vast majority (about 96% below $0.1$), indicating weak conformity penalties for departing from the most recent group norm. (c) The combination, high $d_{i}$ but weak $\phi_{i}$, implies a fragile high state: without an early cooperative majority to anchor behavior, contributions tend to relax downward. This micro pattern mirrors the reduced–form markers: an early tipping threshold near 5.75 L, a village critical mass around 60%, and asymmetric persistence with rarer $L\!\to\!H$ escapes than $H\!\to\!L$ drops.
@@ Appendix I Type classification: Free Riders, Conditional cooperators, Triangles and Other
Using player–level empirical-Bayes slopes from a random-slope learning model and an evidence-based curvature screen, we classify updating rules into Conditional cooperator, Free rider, Triangle, and Other. The aggregate shares are: 48.4% conditional cooperators, 8.5% free riders, 3.1% triangle, and 40.0% other (Table I.1). Unsupervised Ward–D2 clustering on the full ten-round trajectories confirms a two-regime structure (average silhouette =0.360=0.360 at $k=2$; Table 2). The confusion matrix shows that cluster C2 is predominantly Low (LL =641/848=641/848), whereas cluster C1 contains most High finishers and movers (HH =1,038=1{,}038; HL =306=306; LH =171=171) (Table 4). Free riders concentrate in the Low-path cluster (25.9% of C2; essentially absent in C1), producing quasi-separation in Low-path logits (Table I.2, Table I.3). In a mixed-effects specification with type interactions, conditional and triangle types have significantly positive responses to peers’ lag ($\beta_{\bar{c}_{-i}^{t-1}\times\text{Cond}}=0.347,\ p<10^{-12}$; $\beta_{\bar{c}_{-i}^{t-1}\times\text{Tri}}=0.355,\ p<10^{-4}$), while the Other group responds negatively ($\beta=-0.359,\ p<10^{-13}$); the free-rider baseline shows no meaningful peer response (Table I.4). Types are fairly stable over time (early vs. late agreement $75.9\%$; Table I.5), and triangle players’ response peaks at about $0.38$ within-player SD above their typical peer environment (IQR $\approx 0.18$$0.55$; Table I.6). At the village level, a higher free-rider share predicts a larger Low-path fraction (OLS coefficient $0.643\pm 0.066$, $p<0.001$; Table I.7), aligning the micro composition with macro persistence.
@@
Figure I.1 plots $c_{it}$ against $\bar{c}_{-i,t-1}$ with binned means (95% CIs) and type-specific loess lines. The distribution of types is: Conditional cooperators $48.4\%$, Free riders $8.5\%$, Triangle $3.1\%$, Other $40.0\%$ (Table I.1). Conditional and triangle types exhibit robust positive conditional responses, whereas free riders’ slope is near zero and the residual “other” group responds weakly or negatively on average (Table I.4). Using full-trajectory Ward–D2 clustering, the two regimes are well separated (average silhouette $0.360$ at $k=2$; Table 2): C1 concentrates High finishers and movers (HH =1,038=1{,}038, HL =306=306, LH =171=171) while C2 is predominantly Low (LL =641=641) (Table 4). Free riders are concentrated in C2 (25.9% of C2; essentially none in C1) (Table I.2). Types are reasonably stable over time (agreement between rounds 1–5 and 6–10 $75.9\%$; Table I.5). For triangle players the estimated peak occurs at $\hat{x}^{\star}\approx 0.38$ within-player SD above their typical peer environment (IQR $\approx 0.18$$0.55$; Table I.6). At the group level, a larger free-rider share predicts more Low-path membership (OLS coefficient $0.643\pm 0.066$, $p<0.001$; Table I.7).
@@
We resample villages with replacement, re‑cluster the full 10‑round trajectories with Ward–D2 at $k=2$, and record the share assigned to the Low‑path cluster. The distribution centers slightly above one‑half with substantial mass between $\approx 0.3$ and $\approx 0.65$, confirming that the Low basin is not a small‑sample artifact.
Figure L.2: Each draw resamples villages with replacement, reclusters full 10‑round trajectories using Ward–D2 on round‑wise zz–scores ($k=2$), and tags the cluster with the lower mean as the Low path. The distribution is centered a bit above one-half with substantial mass in the 0.300.300.650.65 range, indicating that a sizeable Low basin is not a small‑sample artifact.

### v2

@@ Appendix B Additional data details - Tables & Figures
This appendix documents the sample composition and key distributional facts that underlie the main analyses. We report continuous/count variables in Table B.1, binary indicators in Table B.3, and multi–category variables in Table B.3. Figure B.1 summarizes how the mass of contributions moves across ranges over the ten rounds, and Table B.4 provides corresponding headcounts by round and type. All the percentages are balanced and indicative to the real percentages of the population in Honduran Villages (see Oles2024Maternal; airoldi2024induction; papamichalis2025educational for comparison).
@@ Appendix E IV theory & diagnostics
Each round $t$, player $i$ in a group of size $N=5$ chooses $c_{it}\in[0,12]$. Instantaneous utility combines material payoffs, altruism, and aversion to being at the norm:
$\pi_{it}=\frac{b}{N}\!\left(c_{it}+(N-1)\bar{c}_{-i,t}\right)-\kappa c_{it}+d_{i}c_{it}^{\alpha}-h_{i}\exp\!\left\{k_{\text{norm}}\big(c_{it}-\bar{c}_{-i,t-1}\big)^{2}\right\},$
with $\alpha\in(0,1)$, $d_{i}>0$, $h_{i}\in[0,1]$, $k_{\text{norm}}>0$, and $\bar{c}_{-i,t}$ the leave–one–out peer mean. Under $b/N<\kappa<b$, best–reply dynamics admit two locally attracting fixed points, generating the two–basin structure we document.
@@
$Z^{SS}_{i,t-1}\;=\;\sum_{l}s_{g,-i}^{(l)}\,\mu_{t-1}^{(l)},\qquad\text{and its leave–one–village variant}\quad Z^{LOV}_{i,t-1}\;=\;\sum_{l}s_{g,-i}^{(l)}\,\mu_{t-1}^{(-v),(l)},$
where $s_{g,-i}^{(l)}$ is the LOO group share for trait $l$ and $\mu_{t-1}^{(-v),(l)}$ is the cross–village mean at t1t-1 excluding village $v$. Random assignment within villages and the predetermined nature of traits support relevance and exclusion. Because avatars are anonymous, composition can affect $c_{it}$ only via peers’ contributions.
@@ Appendix F Structural back‑out
We summarize behavior using two primitives. First, an altruism weight $d_{i}$ governing the intrinsic marginal benefit from giving (warm–glow). Second, an effective norm–aversion $\phi_{i}$ that penalizes being at the group’s lagged mean. In the first–order condition, only the product of a curvature parameter and a conformity scale is identified, so we collect these in
$\phi_{i}\equiv 2k_{\text{norm}}\,h_{i},$
and estimate $(d_{i},\phi_{i})$ by minimum distance on each player’s within–session choices using the leave–one–out lagged peer mean (units: Lempiras). [^13] All peer means exclude the focal player. Variables are in lempiras (0–12). Identification is partial: only the product $\phi_{i}=2k_{\text{norm}}h_{i}$ is point–identified from the quadratic deviation cost; the mapping to $(k,h_{i})$ is conventional and used for scaling only. We fix $\alpha=\tfrac{1}{2}$ in the warm–glow term and show sensitivity to $\alpha\in\{0.3,0.7\}$ in SI. Data handling follows the main text: rounds 221010 for lags; a small ridge ($10^{-3}$) when $c_{it}=0$; and 99th–percentile winsorization for the back–out residual.
Three facts emerge. (a) The distribution of $d_{i}$ is broad (median $\approx 2.8$; IQR 1.91.93.43.4); plugging these into the individual best–reply delivers an implied high fixed point at the 12 L cap for essentially all players, i.e., altruism is in principle sufficient to support high contributions. (b) The inferred norm–pull $\phi_{i}$ is near zero for the vast majority (about 96% below $0.1$), indicating weak aversion to matching the most recent group norm. (c) The combination, high $d_{i}$ but weak $\phi_{i}$, implies a fragile high state: without an early cooperative majority to anchor behavior, contributions tend to relax downward. This micro pattern mirrors the reduced–form markers: an early tipping threshold near 5.75 L, a village critical mass around 60%, and asymmetric persistence with rarer $L\!\to\!H$ escapes than $H\!\to\!L$ drops.
@@ Appendix H Illustrative village examples
Figure H.2: Village 4, where most women (pink) are selected from men (blue)
@@ Appendix I Type classification: Free Riders, Conditional cooperators, Triangles and Other
Using player–level empirical-Bayes slopes from a random-slope learning model and an evidence-based curvature screen, we classify updating rules into Conditional cooperator, Free rider, Triangle, and Other. The aggregate shares are: 48.4% conditional cooperators, 8.5% free riders, 3.1% triangle, and 40.0% other (Table I.1). Unsupervised Ward–D2 clustering on the full ten-round trajectories confirms a two-regime structure (average silhouette =0.360=0.360 at two clusters (KK=2); Table 2). The confusion matrix shows that cluster C2 is predominantly Low (LL =641/848=641/848), whereas cluster C1 contains most High finishers and movers (HH =1,038=1{,}038; HL =306=306; LH =171=171) (Table 4). Free riders concentrate in the Low-path cluster (25.9% of C2; essentially absent in C1), producing quasi-separation in Low-path logits (Table I.2, Table I.3). In a mixed-effects specification with type interactions, conditional and triangle types have significantly positive responses to peers’ lag ($\beta_{\bar{c}_{-i}^{t-1}\times\text{Cond}}=0.347,\ p<10^{-12}$; $\beta_{\bar{c}_{-i}^{t-1}\times\text{Tri}}=0.355,\ p<10^{-4}$), while the Other group responds negatively ($\beta=-0.359,\ p<10^{-13}$); the free-rider baseline shows no meaningful peer response (Table I.4). Types are fairly stable over time (early vs. late agreement $75.9\%$; Table I.5), and triangle players’ response peaks at about $0.38$ within-player SD above their typical peer environment (IQR $\approx 0.18$$0.55$; Table I.6). At the village level, a higher free-rider share predicts a larger Low-path fraction (OLS coefficient $0.643\pm 0.066$, $p<0.001$; Table I.7), aligning the micro composition with macro persistence.
@@
Figure I.1 plots $c_{it}$ against $\bar{c}_{-i,t-1}$ with binned means (95% CIs) and type-specific loess lines. The distribution of types is: Conditional cooperators $48.4\%$, Free riders $8.5\%$, Triangle $3.1\%$, Other $40.0\%$ (Table I.1). Conditional and triangle types exhibit robust positive conditional responses, whereas free riders’ slope is near zero and the residual “other” group responds weakly or negatively on average (Table I.4). Using full-trajectory Ward–D2 clustering, the two regimes are well separated (average silhouette $0.360$ at two clusters (KK=2); Table 2): C1 concentrates High finishers and movers (HH =1,038=1{,}038, HL =306=306, LH =171=171) while C2 is predominantly Low (LL =641=641) (Table 4). Free riders are concentrated in C2 (25.9% of C2; essentially none in C1) (Table I.2). Types are reasonably stable over time (agreement between rounds 1–5 and 6–10 $75.9\%$; Table I.5). For triangle players the estimated peak occurs at $\hat{x}^{\star}\approx 0.38$ within-player SD above their typical peer environment (IQR $\approx 0.18$$0.55$; Table I.6). At the group level, a larger free-rider share predicts more Low-path membership (OLS coefficient $0.643\pm 0.066$, $p<0.001$; Table I.7).
@@
We resample villages with replacement, re‑cluster the full 10‑round trajectories with Ward–D2 at two clusters KK=2), and record the share assigned to the Low‑path cluster. The distribution centers slightly above one‑half with substantial mass between $\approx 0.3$ and $\approx 0.65$, confirming that the Low basin is not a small‑sample artifact.
Figure L.2: Each draw resamples villages with replacement, reclusters full 10‑round trajectories using Ward–D2 on round‑wise zz–scores (at two clusters KK=2), and tags the cluster with the lower mean as the Low path. The distribution is centered a bit above one-half with substantial mass in the 0.300.300.650.65 range, indicating that a sizeable Low basin is not a small‑sample artifact.